cost POOL

Alonzo Inc. manufactures two products: car wheels and truck wheels. To determine the amount of overhead to assign to each product line, the controller, YuYu Ortega, has developed the following information:



Car
Truck
Estimated wheels produced
43,000
20,000
Direct labour hours per wheel
2
4


Total estimated overhead costs for the two product lines are 830,000.


A) Calculate the overhead cost assigned to the car wheels and truck wheels, assuming that direct labour hours are used to allocate overhead costs.

Car wheels
enter a dollar amount
Truck wheels
enter a dollar amount
Total overhead
enter a total amount


B) YuYu is not satisfied with the traditional method of allocating overhead because he believes that most of the overhead costs relate to the truck wheel product line because of its complexity. He therefore develops the following three activity cost pools and related cost drivers to better understand these costs:

Activity Cost Pools
Estimated Use of Cost Drivers
Estimated Overhead Costs
Setting up machines
800set-ups
408,000
Assembling
37,000labour hours
370,000
Inspection
104inspections
52,000


Calculate the activity-based overhead rates for these three cost pools.

Cost Pool
Rate
Setting up machines
enter a dollar amountper set-up
Assembling
enter a dollar amountper labour hr
Inspection
enter a dollar amountper inspection
Answer #1
A

WORKINGANS
TOTAL LABOUR HRS (43,000*2)+(20,000*4)=166,000
CAR WHEELS(830,000/166,000)*(43,000*2)=430,000
TRUCK WHEELS(830,000/166,000)*(20,000*4)=400,000
TOTAL OVERHEAD
(430,000+400,000)=830,000

B


ANS
WORKING
COST POOLRATE

SETTING UP MACHINES
510PER SET-UP40,800/800
ASSEMBLING
10PER LABOUR HR370,000/37,000
INSPECTION
500PER INSPECTION52,000/104


answered by: lestat